# The strong convergence phenomenon

### v1

@@ 5.4.1. Tensor models
Here $P$ is any noncommutative polynomial with matrix coefficients of dimension $D$, $\ell$ is an absolute constant, $C$ is a constant that depends only on the degree of $P$, and $\nu_{0},\nu_{1}$ are Schwartz distributions that are supported in $[-\|P(\bm{u},\bm{u}^{*})\|,\|P(\bm{u},\bm{u}^{*})\|]$. If we choose a test function $h$ that vanishes in the latter interval, we obtain
@@
where all $\nu_{i}$ are Schwartz distributions. The new ingredient that now arises is that we must show that the support of each $\nu_{i}$ is included in $[-\|P(\bm{u},\bm{u}^{*})\|,\|P(\bm{u},\bm{u}^{*})\|]$. Surprisingly, a very simple technique that is developed in [33] (see also [110, 109]) shows that this property follows automatically in the present setting from concentration of measure. This yields strong convergence for $D=o(N^{m})$ for any $m\in\mathbb{N}$. Reaching $D=e^{o(N)}$ is harder and requires several additional ideas.
@@ 5.4.2. Some ideas behind the reduction
with probability $1-o(1)$ as $N\to\infty$. The crux of the proof is now to show that $h(M:L(\mathbf{F}_{r}))=0$ implies “microstate collapse”: with high probability, there is a unitary matrix $V$ so that $H_{i}^{N}\approx V\tilde{H}_{i}^{N}V^{*}$ for all $i$. Thus (5.5) contradicts (5.4), and we have achieved the desired conclusion.
@@
Since $\bm{\tilde{H}}^{N}$ is an independent copy of $\bm{H}^{N}$ and thus satisfies the same property, it follows that $d^{\rm orb}(\bm{H}^{N},\bm{\tilde{H}}^{N})=o(1)$ with probability $1-o(1)$.
@@ 5.5.1. Harmonic maps
A critical point of the energy is called a harmonic map. If $f$ is weakly conformal (i.e., conformal away from branch points), then $\mathrm{E}(f)$ coincides with the area of the surface $f(X)$ in $M$. Thus a weakly conformal map $f$ is harmonic if and only if $f(X)$ is a minimal surface in $M$. See, e.g., [102, §4.2.1].
This viewpoint yields a variational method for constructing minimal surfaces. Clearly any minimizer of the energy is, by definition, a harmonic map. In general, such a map is not guaranteed to be weakly conformal. However, this will be the case if we take $X$ to be a surface with a unique conformal class—the thrice punctured sphere—and then a minimizer of $\mathrm{E}(f)$ automatically defines a minimal surface $f(X)$ in $M$. We will make this choice of $X$ from now on. [^22] More generally, one obtains minimal surfaces by minimizing the energy both with respect to the map $f$ and with respect to the conformal class of $X$; see [102, Theorem 4.8.6].
The construction in [122] uses a variant of the variational method which produces minimal surfaces that have many symmetries. Let us write $\𝑋\ΓX=\Gamma\backslash\mathbb{H}$, and consider a unitary representation $\pi_{N}:\Gamma\to U(N)$ with finite range $|\pi_{N}(\Gamma)|<\infty$ which we view as acting on the unit sphere $\mathbb{S}^{2N-1}$ of $\mathbb{C}^{N}$ with its standard Euclidean metric. The following variational problem is considered in [122]:
@@
where $F$ is the fundamental domain of the action of $\Gamma$ on $\mathbb{H}$. To interpret this variational problem, note that a $\pi_{N}$-equivariant map $f:\mathbb{H}\to\mathbb{S}^{2N-1}$ can be identified with a map $f:X^{N}\to\mathbb{S}^{2N-1}$ on the surface $X^{N}=\Gamma_{N}\backslash\mathbb{H}$, where [^23] As $\pi_{N}$ has finite range, $\Gamma_{N}$ is a finite index subgroup of $\Gamma$ and thus $X^{N}$ is a finite cover of $X$. This construction of covering spaces is different than the one considered in section 5.2.
@@
Since a minimizer $f_{N}$ in $\mathrm{E}(X,\pi_{N})$ minimizes the Dirichlet energy, it defines [^24] Even though $X^{N}$ has punctures, taking the closure of $f_{N}(X^{N})$ yields a closed surface. This is a nontrivial property of harmonic maps, see [102, §4.6.4]. a minimal surface $f_{N}(X^{N})$ in $\mathbb{S}^{2N-1}$ that has many symmetries (it contains many rotated copies of the image $f_{N}(F)$ of the fundamental domain).
@@ 5.5.2. An infinite-dimensional model
More precisely, we have the following [122, Corollary 2.4]. Here two unitary representations $\rho_{1}:\Gamma\to B(H_{1})$ and $\rho_{2}:\Gamma\to B(H_{2})$ are said to be weakly equivalent if any matrix element of $\rho_{1}$ can be approximated uniformly on compacts by finite linear combinations of matrix elements of $\rho_{2}$, and vice versa.
@@
Let $\rho:\Gamma\to B(H)$ be a unitary representation of $\Gamma$ that is weakly equivalent to the regular representation $\lambda_{\Gamma}$. Then any $\rho$-equivariant energy-minimizing map $\varphi:\mathbb{H}\to\mathbb{S}^{\infty}$ must satisfy $\varphi^{*}g_{\mathbb{S}^{\infty}}=\frac{1}{8}g_{\rm hyp}$, where $g_{\rm hyp}$ denotes the hyperbolic metric on $\mathbb{H}$ (so $\frac{1}{8}g_{\rm hyp}$ is the metric on $\mathbb{H}$ with constant curvature $-8$).
@@
Theorem 5.14 explains the utility of constructing equivariant minimal surfaces: if we choose the sequence of representations $\pi_{N}$ in such a way that the limiting representation $\pi_{\infty}$ is weakly equivalent to the regular representation, then this will automatically imply that the metrics $f_{N}^{*}g_{\mathbb{S}^{2N-1}}$ on the minimal surfaces converge to the metric $f_{\infty}^{*}g_{\mathbb{S}^{\infty}}=\frac{1}{8}g_{\rm hyp}$ with constant curvature $-8$.
@@ 3. $\bullet$
On the other hand, not every discrete group admits a strongly convergent model: there cannot be a model of random permutation matrices that strongly converges to the regular representation of $\mathbf{F}_{2}\times\mathbf{F}_{2}\times\mathbf{F}_{2}$ (a very special case of a right-angled Artin group) [88, Proposition 2.7], or a model of random unitary matrices that converges strongly to the regular representation of $\mathrm{SL}_{d}(\mathbb{Z})$ with $d\geq 4$ [89]. Thus existence of strongly convergent models cannot be taken for granted.
@@
The most natural random matrix model of this group is obtained by sampling $4$-tuples of random permutation matrices $U_{1}^{N},U_{2}^{N},U_{3}^{N},U_{4}^{N}$ uniformly at random from the set of such matrices that satisfy $[U_{1}^{N},U_{2}^{N}][U_{3}^{N},U_{4}^{N}]=\mathbf{1}$. This constraint introduces complicated dependencies, which causes the model to behave very differently than independent random permutation matrices. For example, unlike in the setting of section 3, the expected traces of monomials of these matrices are not even analytic, let alone rational, as a function of $\frac{1}{N}$.
For surface groups, one can use the representation theory of $\mathbf{S}_{N}$ to analyze this model; in particular, this enabled Magee–Naud–Puder [91] to show that its spectral statistics admit an asymptotic expansion in $\frac{1}{N}$. The proof of strong convergence of this model in [92] is made possible by an extension of the polynomial method to models that admit “good” asymptotic expansions.
@@
This is the fundamental group of the Gieseking manifold, which is obtained by gluing the sides of a tetrahedron [94, §V.2]. Whether sampling uniformly from the set of permutation matrices $U_{1}^{N},U_{2}^{N}$ with $(U_{1}^{N})^{2}(U_{2}^{N})^{2}=U_{2}^{N}U_{1}^{N}$ yields a strongly convergent model is not known. Such questions are of considerable interest, since they provide a route to extending Buser’s conjecture to higher dimensions.
@@ 6.2. Random Cayley graphs
Let $\bm{\sigma}=(\sigma_{1},\ldots,\sigma_{r})$ be i.i.d. uniform random elements of $\mathbf{S}_{N}$, and let $\pi_{N}$ be an irreducible representation of $\mathbf{S}_{N}$. The results in section 5.3 show that the random matrices $\pi_{N}(\bm{\sigma})$ strongly converge to the regular representation $\lambda(\bm{g})$ of the free generators $\bm{g}=(g_{1},\ldots,g_{r})$ of $\mathbf{F}_{r}$ for any sequence of irreducible representations with $1<\dim(\pi_{N})\leq\exp(N^{\frac{1}{12}-\delta})$. What happens beyond this regime is a mystery: it may even be the case that strong convergence holds for any irreducible representations with $\dim(\pi_{N})\to\infty$.
@@
Note that random Cayley graphs of $\mathbf{S}_{N}$ cannot have an optimal spectral gap with probability $1-o(1)$, as there is a nontrivial $1$-dimensional representation (the sign representation). The latter produces a single eigenvalue that is distributed as twice the sum of $r$ independent Bernoulli variables, which exceeds the lower bound of Lemma 1.2 with constant probability. Thus we interpret the optimal spectral gap question to mean whether all eigenvalues, except those coming from the trivial and sign representations, meet the lower bound with probability $1-o(1)$. That this is the case is a well known conjecture, see, e.g., [119, Conjecture 1.6]. However, to date it has not even been shown that such graphs are expanders, i.e., that their nontrivial eigenvalues are bounded away from the trivial eigenvalue as $N\to\infty$; nor is there any known construction of Cayley graphs of $\mathbf{S}_{N}$ that achieve an optimal spectral gap. The only known result, due to Kassabov [79], is that there exists a choice of generators for which the Cayley graph is an expander.
@@ 6.4. Deterministic constructions
To date, all known instances of the strong convergence phenomenon require random constructions (except in amenable situations, cf. [88, §2.1]). This is in contrast to the setting of graphs with an optimal spectral gap, for which explicit deterministic constructions exist and even predate the understanding of random graphs [87, 97]. It remains a major challenge to achieve strong convergence by a deterministic construction.
A potential candidate arises from the celebrated works of Lubotzky–Phillips–Sarnak [87] and Margulis [97], who show that the Cayley graph of $\mathrm{PSL}_{2}(\mathbb{F}_{q})$ defined by a certain explicit deterministic choice of generators has an optimal spectral gap. We may therefore ask, by extension, whether the matrices obtained by applying the regular representation of $\mathrm{PSL}_{2}(\mathbb{F}_{q})$ to these generators converge strongly to the regular representation of the free generators of $\mathbf{F}_{r}$ (cf. section 6.2). This question was raised by Voiculescu [127, p. 146] in an early paper that motivated the development of strong convergence of random matrices by Haagerup and Thorbjørnsen. However, the deterministic question remains open, and the methods of [87, 97] appear to be powerless for addressing this question.
@@ 6.5. Ramanujan constructions
The interest in Ramanujan graphs stems in part from an analogy with number theory: the Ramanujan property of a graph is equivalent to the validity of the Riemann hypothesis for its Ihara zeta function [124, Theorem 7.4]. In the setting of hyperbolic surfaces, the analogous “Ramanujan property” that a hyperbolic surface $X$ has $\lambda_{1}(X)\geq\frac{1}{4}$ is equivalent to the validity of the Riemann hypothesis for its Selberg zeta function [104, §6]. An important conjecture of Selberg [120] predicts that a specific family of hyperbolic surfaces has this property. However, no such surfaces have yet been proved to exist. The results in section 5.2 therefore provide additional motivation for studying “Ramanujan forms” of strong convergence.
@@ 6.6. The optimal dimension of matrix coefficients
The strong convergence problem for polynomials of $N$-dimensional random matrices with matrix coefficients of dimension $D_{N}\to\infty$ was discussed in section 5.4.1 in the context of the Peterson-Thom conjecture. While only the case $D_{N}=N$ is needed for that purpose, the optimal range of $D_{N}$ for which strong convergence holds remains open: for both Gaussian and Haar distributed matrices, it is known that strong convergence holds when $D_{N}=e^{o(N)}$ and can fail when $D_{N}\geq e^{CN^{2}}$ [33]. Understanding what lies in between is related to questions in operator space theory [116, §4].
@@ 6.9. Complex eigenvalues
It is not even entirely obvious at first sight how the complex eigenvalue distribution of a non-self-adjoint operator in a $C^{*}$-probability space should be defined. The natural object of this kind, whose definition reduces to the complex eigenvalue distribution in the case of matrices, is called the Brown measure [99, Chapter 11]. It is tempting to conjecture that if a family of random matrices $\bm{X}^{N}$ strongly converges to a family of limiting operators $\bm{x}$, then the empirical distribution of the complex eigenvalues of any noncommutative polynomial $P(\bm{X}^{N},\bm{X}^{N*})$ should converge to $P(\bm{x},\bm{x}^{*})$. To date, this has only been proved in the special case of quadratic polynomials of independent complex Ginibre matrices [44].

### v2

@@ 5.4.1. Tensor models
Here $P$ is any noncommutative polynomial with matrix coefficients of dimension $D$, $\ell$ is an absolute constant, $C$ is a constant that depends only on the degree of $P$, and $\nu_{0},\nu_{1}$ are Schwartz distributions that are supported in $\|P(\bm{u},\bm{u}^{*})\|,\|P(\bm{u},\bm{u}^{*})\|]$. If we choose a test function $h$ that vanishes in the latter interval, we obtain
@@
where all $\nu_{i}$ are Schwartz distributions. The new ingredient that now arises is that we must show that the support of each $\nu_{i}$ is included in $\|P(\bm{u},\bm{u}^{*})\|,\|P(\bm{u},\bm{u}^{*})\|]$. Surprisingly, a very simple technique that is developed in [33] (see also [111, 113]) shows that this property follows automatically in the present setting from concentration of measure. This yields strong convergence for $D=o(N^{m})$ for any $m\in\mathbb{N}$. Reaching $D=e^{o(N)}$ is harder and requires several additional ideas.
@@ 5.4.2. Some ideas behind the reduction
with probability 1o(1)1-o(1) as $N\to\infty$. The crux of the proof is now to show that $h(M:L(\mathbf{F}_{r}))=0$ implies “microstate collapse”: with high probability, there is a unitary matrix $V$ so that $H_{i}^{N}\approx V\tilde{H}_{i}^{N}V^{*}$ for all $i$. Thus (5.5) contradicts (5.4), and we have achieved the desired conclusion.
@@
Since $\bm{\tilde{H}}^{N}$ is an independent copy of $\bm{H}^{N}$ and thus satisfies the same property, it follows that $d^{\rm orb}(\bm{H}^{N},\bm{\tilde{H}}^{N})=o(1)$ with probability 1o(1)1-o(1).
@@ 5.5.1. Harmonic maps
A critical point of the energy is called a harmonic map. If $f$ is weakly conformal (i.e., conformal away from branch points), then $\mathrm{E}(f)$ coincides with the area of the surface $f(X)$ in $M$. Thus a weakly conformal map $f$ is harmonic if and only if $f(X)$ is a minimal surface in $M$. See, e.g., [104, §4.2.1].
This viewpoint yields a variational method for constructing minimal surfaces. Clearly any minimizer of the energy is, by definition, a harmonic map. In general, such a map is not guaranteed to be weakly conformal. However, this will be the case if we take $X$ to be a surface with a unique conformal class—the thrice punctured sphere—and then a minimizer of $\mathrm{E}(f)$ automatically defines a minimal surface $f(X)$ in $M$. We will make this choice of $X$ from now on. [^22] More generally, one obtains minimal surfaces by minimizing the energy both with respect to the map $f$ and with respect to the conformal class of $X$; see [104, Theorem 4.8.6].
@@
$\mathrm{E}(X,\pi_{N})=\inf\bigg\{\frac{1}{2}\int_{F}|df|^{2};~f:\mathbb{H}\to\mathbb{S}^{2N-1}\text{ is }\pi_{N}\text{-equivariant}\bigg\},$
[…]
@@
Since a minimizer $f_{N}$ in $\mathrm{E}(X,\pi_{N})$ minimizes the Dirichlet energy, it defines [^24] Even though $X^{N}$ has punctures, taking the closure of $f_{N}(X^{N})$ yields a closed surface. This is a nontrivial property of harmonic maps, see [104, §4.6.4]. a minimal surface $f_{N}(X^{N})$ in $\mathbb{S}^{2N-1}$ that has many symmetries (it contains many rotated copies of the image $f_{N}(F)$ of the fundamental domain).
@@ 5.5.2. An infinite-dimensional model
More precisely, we have the following [124, Corollary 2.4]. Here two unitary representations $\rho_{1}:\Gamma\to B(H_{1})$ and $\rho_{2}:\Gamma\to B(H_{2})$ are said to be weakly equivalent if any matrix element of $\rho_{1}$ can be approximated uniformly on compacts by finite linear combinations of matrix elements of $\rho_{2}$, and vice versa.
@@
Let $\rho:\Gamma\to B(H)$ be a unitary representation of $\Gamma$ that is weakly equivalent to the regular representation $\lambda_{\Gamma}$. Then any $\rho$-equivariant energy-minimizing map $\varphi:\mathbb{H}\to\mathbb{S}^{\infty}$ must satisfy $\varphi^{*}g_{\mathbb{S}^{\infty}}=\frac{1}{8}g_{\rm hyp}$, where $g_{\rm hyp}$ denotes the hyperbolic metric on $\mathbb{H}$ (so $\frac{1}{8}g_{\rm hyp}$ is the metric on $\mathbb{H}$ with constant curvature 8-8).
@@
Theorem 5.14 explains the utility of constructing equivariant minimal surfaces: if we choose the sequence of representations $\pi_{N}$ in such a way that the limiting representation $\pi_{\infty}$ is weakly equivalent to the regular representation, then this will automatically imply that the metrics $f_{N}^{*}g_{\mathbb{S}^{2N-1}}$ on the minimal surfaces converge to the metric $f_{\infty}^{*}g_{\mathbb{S}^{\infty}}=\frac{1}{8}g_{\rm hyp}$ with constant curvature 8-8.
@@ 3. $\bullet$
On the other hand, not every discrete group admits a strongly convergent model: there cannot be a model of random permutation matrices that strongly converges to the regular representation of $\mathbf{F}_{2}\times\mathbf{F}_{2}\times\mathbf{F}_{2}$ (a very special case of a right-angled Artin group) [89, Proposition 2.7], or a model of random unitary matrices that converges strongly to the regular representation of $\mathrm{SL}_{d}(\mathbb{Z})$ with $d\geq 4$ [90]. Thus existence of strongly convergent models cannot be taken for granted.
@@
The most natural random matrix model of this group is obtained by sampling $4$-tuples of random permutation matrices $U_{1}^{N},U_{2}^{N},U_{3}^{N},U_{4}^{N}$ uniformly at random from the set of such matrices that satisfy $U_{1}^{N},U_{2}^{N}][U_{3}^{N},U_{4}^{N}]=\mathbf{1}$. This constraint introduces complicated dependencies, which causes the model to behave very differently than independent random permutation matrices. For example, unlike in the setting of section 3, the expected traces of monomials of these matrices are not even analytic, let alone rational, as a function of $\frac{1}{N}$.
For surface groups, one can use the representation theory of $\mathbf{S}_{N}$ to analyze this model; in particular, this enabled Magee and Puder [93] to show that its spectral statistics admit an asymptotic expansion in $\frac{1}{N}$. The proof of strong convergence of this model in [94] is made possible by an extension of the polynomial method to models that admit “good” asymptotic expansions.
@@
This is the fundamental group of the Gieseking manifold, which is obtained by gluing the sides of a tetrahedron [96, §V.2]. Whether sampling uniformly from the set of permutation matrices $U_{1}^{N},U_{2}^{N}$ with $(U_{1}^{N})^{2}(U_{2}^{N})^{2}=U_{2}^{N}U_{1}^{N}$ yields a strongly convergent model is not known. Such questions are of considerable interest, since they provide a route to extending Buser’s conjecture to higher dimensions.
@@ 6.2. Random Cayley graphs
Let $\bm{\sigma}=(\sigma_{1},\ldots,\sigma_{r})$ be i.i.d. uniform random elements of $\mathbf{S}_{N}$, and let $\pi_{N}$ be an irreducible representation of $\mathbf{S}_{N}$. The results in section 5.3 show that the random matrices $\pi_{N}(\bm{\sigma})$ strongly converge to the regular representation $\lambda(\bm{g})$ of the free generators $\bm{g}=(g_{1},\ldots,g_{r})$ of $\mathbf{F}_{r}$ for any sequence of irreducible representations with $1<\dim(\pi_{N})\leq\exp(N^{\frac{1}{20}-\delta})$. What happens beyond this regime is a mystery: it may even be the case that strong convergence holds for any irreducible representations with $\dim(\pi_{N})\to\infty$.
@@
Note that random Cayley graphs of $\mathbf{S}_{N}$ cannot have an optimal spectral gap with probability 1o(1)1-o(1), as there is a nontrivial $1$-dimensional representation (the sign representation). The latter produces a single eigenvalue that is distributed as twice the sum of $r$ independent Bernoulli variables, which exceeds the lower bound of Lemma 1.2 with constant probability. Thus we interpret the optimal spectral gap question to mean whether all eigenvalues, except those coming from the trivial and sign representations, meet the lower bound with probability 1o(1)1-o(1). That this is the case is a well known conjecture, see, e.g., [121, Conjecture 1.6]. However, to date it has not even been shown that such graphs are expanders, i.e., that their nontrivial eigenvalues are bounded away from the trivial eigenvalue as $N\to\infty$; nor is there any known construction of Cayley graphs of $\mathbf{S}_{N}$ that achieve an optimal spectral gap. The only known result, due to Kassabov [80], is that there exists a choice of generators for which the Cayley graph is an expander.
@@ 6.4. Deterministic constructions
To date, all known instances of the strong convergence phenomenon require random constructions (except in amenable situations, cf. [89, §2.1]). This is in contrast to the setting of graphs with an optimal spectral gap, for which explicit deterministic constructions exist and even predate the understanding of random graphs [88, 99]. It remains a major challenge to achieve strong convergence by a deterministic construction.
A potential candidate arises from the celebrated works of Lubotzky–Phillips–Sarnak [88] and Margulis [99], who show that the Cayley graph of $\mathrm{PSL}_{2}(\mathbb{F}_{q})$ defined by a certain explicit deterministic choice of generators has an optimal spectral gap. We may therefore ask, by extension, whether the matrices obtained by applying the regular representation of $\mathrm{PSL}_{2}(\mathbb{F}_{q})$ to these generators converge strongly to the regular representation of the free generators of $\mathbf{F}_{r}$ (cf. section 6.2). This question was raised by Voiculescu [129, p. 146] in an early paper that motivated the development of strong convergence of random matrices by Haagerup and Thorbjørnsen. However, the deterministic question remains open, and the methods of [88, 99] appear to be powerless for addressing this question.
@@ 6.5. Ramanujan constructions
The interest in Ramanujan graphs stems in part from an analogy with number theory: the Ramanujan property of a graph is equivalent to the validity of the Riemann hypothesis for its Ihara zeta function [126, Theorem 7.4]. In the setting of hyperbolic surfaces, the analogous “Ramanujan property” that a hyperbolic surface $X$ has $\lambda_{1}(X)\geq\frac{1}{4}$ is equivalent to the validity of the Riemann hypothesis for its Selberg zeta function [106, §6]. An important conjecture of Selberg [122] predicts that a specific family of hyperbolic surfaces has this property. However, no such surfaces have yet been proved to exist. The results in section 5.2 therefore provide additional motivation for studying “Ramanujan forms” of strong convergence.
@@ 6.6. The optimal dimension of matrix coefficients
The strong convergence problem for polynomials of $N$-dimensional random matrices with matrix coefficients of dimension $D_{N}\to\infty$ was discussed in section 5.4.1 in the context of the Peterson-Thom conjecture. While only the case $D_{N}=N$ is needed for that purpose, the optimal range of $D_{N}$ for which strong convergence holds remains open: for both Gaussian and Haar distributed matrices, it is known that strong convergence holds when $D_{N}=e^{o(N)}$ and can fail when $D_{N}\geq e^{CN^{2}}$ [33]. Understanding what lies in between is related to questions in operator space theory [118, §4].
@@ 6.9. Complex eigenvalues
It is not even entirely obvious at first sight how the complex eigenvalue distribution of a non-self-adjoint operator in a $C^{*}$-probability space should be defined. The natural object of this kind, whose definition reduces to the complex eigenvalue distribution in the case of matrices, is called the Brown measure [101, Chapter 11]. It is tempting to conjecture that if a family of random matrices $\bm{X}^{N}$ strongly converges to a family of limiting operators $\bm{x}$, then the empirical distribution of the complex eigenvalues of any noncommutative polynomial $P(\bm{X}^{N},\bm{X}^{N*})$ should converge to $P(\bm{x},\bm{x}^{*})$. To date, this has only been proved in the special case of quadratic polynomials of independent complex Ginibre matrices [44].